A coffee merchant combines coffee costing $5.20/lb with coffee costing $3.00/lb. How many pounds of each should be used to make 40 lbs. of a blend costing $4.00/lb?

Answer by checkley75(3666)

5.2X+3(40-X)=4*40
5.2X+120-3X=160
2.2X=160-120
2.2X=40
X=40/2.2
X=18.18 POUND OF $5.20 COFFEE IS REQUIRED.
40-18.18=21.82 POUNDS OF $3.00 COFFEE IS REQUIRED.
PROOF
5.2*18.18+3*21.82=160
94.54+65.46=160
160=160